If the proportion of an African population that are susceptible to Malaria is 0.449, what is the HbA allele frequency in the population?
Answer #1

Answer

HbA allele frequency = 0.572

Comment:

HbA/HbA (P2) - This homozygous allele combination is susceptible to malaria

HbA/HbS (2PQ) - This homozygous allele combination is resistance to malaria

HbS/HbS (Q2) - This homozygous allele combination will not survive

So, for given question HbA/HbA (P2) = 0.449

So from Hardy Weinberg equation (P+Q) =1

Or Q = 1-P      or, 1-0.449       or, Q= 0.551

Allele frequency of (P) = frequency of HbA/HbA + ½ frequency of HbA/HbS

                                                = 0.449 + ½ (0.449 X 0.551)

                                    0.572